a.      Express the roots of (−14+3i)−2/5 complex number in polar form. 


1
Expert's answer
2021-03-08T19:17:18-0500
"(-14+3i)^{-2}=\dfrac{187}{42025}+\dfrac{84}{42025}i"

"r=\dfrac{1}{205}, \theta=\tan^{-1}(\dfrac{84}{187})"

According to the De Moivre's Formula, all "n"-th roots of a complex number "r(\cos \theta+i\sin \theta)" are given by "\sqrt[n]{r}\bigg(\cos (\dfrac{\theta+2\pi k}{n})+i\sin (\dfrac{\theta+2\pi k}{n})\bigg),"

"k=0, 1, 2, ... n-1"

"k=0:"


"\dfrac{1}{\sqrt[5]{205}}(\cos (\dfrac{\tan^{-1}(\dfrac{84}{187})}{5})+i\sin (\dfrac{\tan^{-1}(\dfrac{84}{187})}{5}))"

"k=1:"

"\dfrac{1}{\sqrt[5]{205}}(\cos (\dfrac{\tan^{-1}(\dfrac{84}{187})+2\pi}{5})+i\sin (\dfrac{\tan^{-1}(\dfrac{84}{187})+2\pi}{5}))"




"k=2:"

"\dfrac{1}{\sqrt[5]{205}}(\cos (\dfrac{\tan^{-1}(\dfrac{84}{187})+4\pi}{5})+i\sin (\dfrac{\tan^{-1}(\dfrac{84}{187})+4\pi}{5}))"



"k=3:"

"\dfrac{1}{\sqrt[5]{205}}(\cos (\dfrac{\tan^{-1}(\dfrac{84}{187})+6\pi}{5})+i\sin (\dfrac{\tan^{-1}(\dfrac{84}{187})+6\pi}{5}))"



"k=4:"

"\dfrac{1}{\sqrt[5]{205}}(\cos (\dfrac{\tan^{-1}(\dfrac{84}{187})+8\pi}{5})+i\sin (\dfrac{\tan^{-1}(\dfrac{84}{187})+8\pi}{5}))"